OpenStudy (anonymous):

Which is an equation of a circle with center (2, -1) that passes through the point (3, 4)? (x - 2)2 + (y + 1)2 = 13 (x + 2)2 + (y - 1)2 = 13 (x + 2)2 + (y - 1)2 = 26 (x - 2)2 + (y + 1)2 = 26

OpenStudy (psymon):

Well, the form is (x-h)^2 + (y-k)^2 = r^2, where the minus does not change. So you be plugging in 2 and -1 into the form of (x-h) and (y-k). and the last part is r^2, which you can find by drawing a triangle using the center of the circle and your second point. If you draw a triangle and use pythagorean theorem, you can find the answer for r^2.

OpenStudy (anonymous):

so r^2 is 26 right?

OpenStudy (psymon):

Looks right to me.

OpenStudy (anonymous):

as mentioned above the general fporm of Equation of circle is \[\Large (x-h)^2+(y-k)^2 =r^2\] here (h,k) is the center of the circle. since center of the circle is given (2,-1) it means h=2 and k=-1 put h=2 and k=-1 in the equation of the circle aftter plugging the values of h and k I am getting \[\Large (x-2)^2+(y-(-1))^2=r^2\] \[\Large (x-2)^2+(y+1)^2=r^2\] we have two possible options which are similar . they are \[\Large (x-2)^2+(y+1)^2=13\] \[\Large (x-2)^2+(y+1)^2=26\] to decide which one is correct put the given point (3,4) in the above two equation. the equation which will satify the point is the Amswer plugging into first equation x=3 and y=4 \[\Large (3-2)^2+(4+1)^2=13\] \[\Large 26 =13\] this is incorrect because 26 is not equal to 13! plug x=3 and y=4 in the second equation \[\Large (3-2)^2+(4+1)^2=26\] \[\Large 26=26\] Yes this is Logical make sense! so the correct Option should be \[\Large (x-2)^2+(y+1)^2=26\]

OpenStudy (anonymous):

@elizabethvilleda